A molecule having a molar mass of 151.0 g/mol crystallizes to a cubic solid where each edge of the unit cell is 714.0 pm. Each unit cell contains 2 molecules. What is the bulk density of the crystal?
1
Expert's answer
2019-09-18T02:49:57-0400
V(2 molecules)=(714*10^-12)^3=3.640*10^-28 m^3
V(1 mole)=6.022*10^23/2*3.640*10^-28*10^6=109.598696978 cm^3
p=151/109.6=1.378 g/cm^3
ANSWER: 1.378 (g/cm^3)